OpenStudy (anonymous):

Prove that: Δ=4.R.r.cos a/2.cos b/2. cos y/2

OpenStudy (anonymous):

\[\frac{ abc }{ 4\Delta } * \frac{ \Delta }{ S } * \sqrt{\frac{ s(s-a) }{ bc }} * \sqrt{\frac{ s(s-b) }{ ac }} * \sqrt{\frac{ s(s-c) }{ ab }}\]

OpenStudy (anonymous):

Using R.H.S

OpenStudy (anonymous):

it is easy just do the multiplication. let the triangle sign be t so rhs= 4[abc/4t][t/s][sqrt(s(s-a)/bc)][sqrt(s(s-b)/ac)][sqrt(s(s-c)/ab)] =[4abct/4ts][(s/abc)sqrt(s(s-a)(s-b)(s-c))] =[abc/s][s/abc]sqrt(s(s-a)(s-b)(s-c)) =sqrt(s(s-a)(s-b)(s-c)) =t hence proved